OpenStudy (redohawk):

Question 1 (Multiple Choice Worth 5 points) (06.04 MC) For the line segment whose endpoints are X (1, 2) and Y (6, 7), find the x value for the point located 1 over 3 the distance from X to Y. 3.5 3.3 2.7 2.4

OpenStudy (redohawk):

@jcpd910 ?

OpenStudy (redohawk):

@Abhisar

OpenStudy (redohawk):

@-BeYourself20- @Mathhelp123344

OpenStudy (redohawk):

Use the following four coordinates to determine which one has a distance of 13 units from point G (13, 2). (7, 2) (9, 5) (6, 7) (1, 7)

OpenStudy (redohawk):

@Kendra1998

OpenStudy (anonymous):

I think its C

OpenStudy (redohawk):

which one?

OpenStudy (redohawk):

@hartnn

OpenStudy (anonymous):

2.7

OpenStudy (redohawk):

reason? or is just a guess?

OpenStudy (anonymous):

I googled it, sorry

OpenStudy (anonymous):

why are you asking me anyway, your smartscore is higher then mine Lol

OpenStudy (redohawk):

do you really believe in numbers? just because i might have more subscribers on youtube then someone else. Does not really mean that my videos are better.

hartnn (hartnn):

for the 1st one, the formula is \(\Large x = \dfrac{mx_2 +nx_1}{m+n}\)

hartnn (hartnn):

where m:n is the ratio. since you need, 1/3rd distance. so ratio will be 1:2 m= 1, n=2

OpenStudy (redohawk):

um alright, which on is M and which is N?

OpenStudy (redohawk):

nvm

hartnn (hartnn):

x1= 1, x2 =6

OpenStudy (redohawk):

so.. x = 1x^2 + 2x^1/1+2?

OpenStudy (redohawk):

i'm stuck :/

hartnn (hartnn):

plug in x1 an d x2 values too

OpenStudy (redohawk):

but what does x equal?

hartnn (hartnn):

x is your answer...you need to find x

hartnn (hartnn):

\(\Large x = \dfrac{mx_2 +nx_1}{m+n}= \dfrac{1\times 6 +2\times 1}{1+2} =\)

OpenStudy (redohawk):

based on what you wrote. it would equal 8/3 8/3 = 2.66666666 which would be 2.7

OpenStudy (redohawk):

@Kendra1998 you were right xD :P

hartnn (hartnn):

yes, but we should know how it is 2.7 :)

OpenStudy (redohawk):

yes, thanks for your help ^^ can you help me with the other one?

hartnn (hartnn):

do you know distance formula ?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (redohawk):

what do we plugin for x?

hartnn (hartnn):

G (13, 2). x1 = 13, y1 = 2

OpenStudy (redohawk):

where did G come from? i feel like an idiot xD

hartnn (hartnn):

take each option , one by one say 1st, 7,2 plug in x2 =7, y2 = 2 see if you get distance = 13

hartnn (hartnn):

from the question

OpenStudy (redohawk):

ahh.

OpenStudy (redohawk):

(x1 - 7)^2 + (y1 - 2)^2 is this right so far @hartnn

hartnn (hartnn):

x1 = 13, y1 = 2

OpenStudy (redohawk):

13 - 7 = 6 6^2 = 36 2-2 = 0 0^2 = 0 so.. 36 + 0 = 36

OpenStudy (redohawk):

so.. D = 36 I am still lost :(

hartnn (hartnn):

and sqrt 36 = 6 so it is not 13 so, it is not 1st choice

hartnn (hartnn):

now try 2nd choice

hartnn (hartnn):

x1 = 13, y1 = 2 x2 =9 y2 =5

OpenStudy (anonymous):

lol @RedoHawk actually google was right but thanks lol

OpenStudy (redohawk):

13-2 = 11 11^2 = 121 9-5 = 4 4^2 = 16 121 + 16 = 137 now what?

OpenStudy (redohawk):

@Kendra1998 lol true :P

OpenStudy (redohawk):

@hartnn what answer do we have to end up with?

hartnn (hartnn):

distance should come to be 13 try this: (1, 7)

hartnn (hartnn):

x1 = 13, y1 = 2 x2 =1 y2 =7

OpenStudy (redohawk):

13-2 = 11 11^2 = 121 1 - 6 = -6 -6^2 = 36 121 + 36 = 157 now what? i don't know why i am doing this. Where did 13 come from? this makes no sense :/

hartnn (hartnn):

you are doing x1-y1 look at the formula again its x1-x2

OpenStudy (redohawk):

13-1 = 12 12^2 = 144 2 - 7 = -5 -5^2 = 25 144 + 25 = 169

OpenStudy (redohawk):

can you just tell me the answer maybe?

hartnn (hartnn):

you got it! \(d= \sqrt{169}=13\) so last choice is the option

hartnn (hartnn):

correct option**

OpenStudy (redohawk):

um okay, thanks ^^ really appreciate it bro :D